Easy Sect. # 2 A club has 8 doctors, 10 enginears, and 12 lawyers, It a person is picked at random, what is the probability that be or she is a lawyer? Easy Sect. 2 engineer? A elub haa 8 doctorn, 10 angineers, and 12 laxyora. It a person in picked at randon, what is the probability that he or she is not an 5-204 It a die ia rolled ance, what is the probability of getting s aeven? Taay Sect. 2 5-105 It a die ia rolled once, what is the prabability ot getting a nusber lens than three? L.?. 2 5-106 ty of greater than or equal to 5? Sect. #2 L.o. 2 5-107 Med Sect. # 2 z.?. If a die i rolled once, hat ?. the probabLl nore than 2 or an even nunber? ity of getting a turbe? 5-108 Med A certain type of ear has a 10t chance of breaking dowm. Out of the next 90 cars, how many wili break down? tho. 5-109Tvo dice are rolled. Waat is the probability of getting a sun of Eaby Sect. 2 1.0, 2 5-110 Two dice are rolled. What is the probability of getting a sun ot s or doubles? Ned Sect. #2 A box containa 9 blue narbles, 11 red marbles, and 3 yellov narbles. It a marble is aelected at randon, what is the probability that it la Kaay Sect. # 2 not zed? L.0.2

needs help with these problems
Answer #1

5.102.

If a person is picked at random, probability that he or she is a lawyer = (total number of lawyers) / (total persons)

= 12 / 30 = 0.4

If a person is picked at random, probability that he or she is a lawyer is 0.4.

5.103

If a person is picked at random, probability that he or she is not an engineer = (total number of non engineers) / (total persons)

= 20 / 30 = 0.67

If a person is picked at random, probability that he or she is not an engineer is 0.67.

5-104

It is not possible of getting seven in rolling single die once. So the probability of getting a seven if a die is rolled once is 0.

5-105

The possible numbers when rolling a die once are {1, 2, 3, 4, 5 and 6}. The number of events less than three is 2.

So the probability of getting number less than three when a die is thrown once is 2 / 6 = 0.33.

If a die is rolled once, the probability of getting a number less than three is 0.33.

5-106

The possible numbers for greater than equal to 5 are 5 and 6 only. So the probability of getting a number greater than or equal to 5 when rolling a die once is 2 / 6 = 0.33.

If a die is rolled once, the probability of getting a number greater than or equal to 5 is 0.33.

5-107

The number of events of getting numbers more than 2 is 4 {3, 4, 5 and 6}

The number of events of getting even numbers is 3 {2, 4 and 6}

In these two events the number 4 and 6 are common. So the event of getting a number greater than 2 or even number is {2, 3, 4, 5 and 6}

The probability of getting number greater than 2 or even number is 5 / 6

5-108

The chance of breakdown of car is 10%. So out of 90 cars, 90 (0.1) = 9 cars are expected to be break down.

5 - 109

When two dice are thrown, the events of getting a sum of 8 are

{(2,6), (3,5), (4,4) (5,3) and (6,2)} and the number of possible outcomes when throwing two dice is 36.

So the probability of getting a sum of eight when throwind two dice are 5 / 36.

5-110

When two dice are thrown, the events of getting a sum of 5 or a doubles are

{(1,1),(1,4),(2,2),(2,3),(3,2),(3,3),(4,1),(4,4),(5,5) and (6,6)} and the number of possible outcomes when throwing two dice is 36.

So the probability of getting a sum of five or a doubles is 10 / 36 = 5 / 18.

5-111

The probabiliy of getting a non red = number of non-red marbles / total number of marbles = 12/23.

Probability that a marble selected at random is non-red is 12 / 23.